# version 0.75

### v4

@@
[…]
position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments and suggestions welcome at nonstochastic.control@gmail.com
Introduction to Online Control
version 0.75
@@ Acknowledgments
This book was a journey that started in Princeton at the Google DeepMind Lab in the summer of 2018. At that time Sham Kakade visited the group, which consisted of the authors as well as Naman Agarwal, Brian Bullins, Xinyi Chen, Cyril Zhang and Yi Zhang, henceforth the gang. The gang worked on various aspects of optimization, reinforcement learning, and dynamical systems. The first paper, in which online nonstochastic control was basically invented, started in room 401 in the computer science building at Princeton University. We remember and cherish the exact moment along with the mathematical writings in white chalk on that blackboard.
@@
The following students were particularly helpful making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna (and many more we wish to announce very soon).
@@
Algebra, Geometry and Calculus
@@ 1.1 What is This Book About?
Control theory is the engineering science of manipulating physical systems to achieve desired functionality. Its history is centuries old, and it spans rich areas from the mathematical sciences. A brief list includes differential equations, operator theory and linear algebra, functional analysis, harmonic analysis, and more. However, it mostly does not concern modern algorithmic theory and computational complexity from computer science.
The advance of the deep learning revolution has brought to light the power and robustness of algorithms that are based on gradient methods. The following quote is attributed to deep learning pioneer Yann Lecun during a Barbados workshop on deep learning theory circa Feb 2020,
@@
This is a deep and difficult question that we fall short of answering. We do, however, propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
@@ 1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century. One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism. A major source of losses in efficiency of earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily result in engine failure. The centrifugal governor was created to solve this problem.
Figure 1.1: A centrifugal governor. Source: Wikipedia
The centrifugal governor consists of a shaft connected to two metal balls that are allowed to rotate. Increased shaft speed causes the balls to spin faster, and the resulting centrifugal force causes the balls to lift. The balls are in turn attached to a pin that controls the amount of steam entering the engine, so as the balls spin faster, the pin lifts and limits the amount of fuel that can enter the system. On the other hand, if the balls spin slower, the pin lowers, allowing more steam to enter the system. This provides a natural, direct and proportional control to the system.
The impact of the centrifugal governor on the steam engine, and in turn on the the industrial revolution, illustrates the importance of the field of control in engineering and the sciences. Since the 18th century our world has changed quite a bit, and the importance of control has only grown.
The challenges of controlling systems has moved from the steam engine to design of aircraft, robotics and autonomous vehicles. The importance of designing efficient and robust methods remains at least as crucial as in the industrial revolution.
@@ 1.3 Formalization and Examples of a Control Problem
With the physical example of the centrifugal governor in mind, we give a mathematical definition of a control problem. The first step is to formally define a dynamical system below.
@@
A dynamical system from starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$ is given by the following equations:
@@
The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$, control input $\mathbf{u}_{t}$ and perturbation $\mathbf{w}_{t}$. The subscript indicates the relevant quantities at time step $t$.
Given a dynamical system, there are several possible objectives we can formulate. We give a generic formulation below that captures many common control formulations, notably the optimal control problem and the robust control problem. These will be addressed in much more detail later in this book. In this problem formulation, the goal is to minimize the long-term horizon cost given by a sequence of cost functions.
@@
At this point, there are many aspects of the generic problem that are left unspecified. These include:
@@
Different theories of control differ by the answers they give to the questions above, and these in turn give rise to different methods.
Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem. We can then reason about the different questions, and motivate the nonstochastc control setting.
@@ 1.3.1 Example: Control of a Medical Ventilator
The ventilator needs to take into account the structure of the lung to determine the optimal pressure to induce. Such structural factors include compliance, or the change in lung volume per unit pressure, and resistance, or the change in pressure per unit flow.
@@
where $p_{t}$ is the observed pressure, $v_{t}$ is the volume of the lung, $u_{t}$ is the input flow (control), and $\Delta_{t}$ is the unit of time measured by the device precision. $C_{0}$, $C_{1}$, and $C_{2}$ are constants that depend on the particular system/lung-ventilator pair.
The goal of ventilator control is to regulate pressure according to some predetermined pattern. As a dynamical system, we can write $v_{t+1}=f(v_{t},u_{t},w_{t})$, $p_{t}=g(v_{t})$, and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$. The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
@@ 1.4 Simple Control Algorithms
The previous section detailed the importance of the control problem on the example of a critical care medical device. In this section we detail methods for controlling a medical ventilator, or indeed any physical device that can be modelled as a dynamical system.
The following methods are simplistic but still widely used. For example, the PID algorithm is still used in the majority of medical ventilators today. In later sections we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
@@ 1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$, that is assumed to contain zero for simplicity. A bang-bang controller chooses the control input at time $t$ as
@@
A bang-bang controller saturates its inputs to extreme values in an attempt to confine the state of the system to a desirable range. For example, in the context of mechanical ventilation, one could choose $[p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$. A bang-bang controller in such a case would let the oxygen in at the maximum possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
@@ 1.4.2 The PID Controller
A special sub-family of linear controllers is that of linear controllers that optimize over a pre-defined basis of historical observations. More precisely,
@@
The PID controller, named after the nature of the three components that constitute it, specifies three coefficients $\alpha,\beta,\gamma$ to generate a control signal. In that sense it is extremely sparse, which from a learning theoretic perspective is a good indication of generalization.
Indeed, the PID controller is the method of choice for numerous applications in engineering, including control of medical ventilators.
@@ 1.5 Classical Theory: Optimal and Robust Control
Classical theories of control differ over the generic control problem formulation 1.2 primarily in one significant aspect: the perturbation model. A rough divide between the two leading theories of optimal control and robust control is as follows.
Optimal control postulates a probabilistic model for the perturbations, which are typically i.i.d. draws from a Gaussian distribution. However, other types of distributions can be considered as well. As such, the control problem can be re-phrased in a more precise stochastic mathematical program,
@@
In contrast, robust control theory plans for the worst case noise from a given set of constraints, denoted $\mathcal{K}$. Thus, the robust control problem can be rephrased as the following mathematical program,
@@
The two formulations have been debated for decades, with pros and cons for each side. Optimal control focuses on the average case and many times yields simpler formulations for optimization, resulting in more efficient methods. Robust control, on the other hand, allows for more difficult noise models and some amount of model misspecification. However, it is too pessimistic for many natural problems.
Nonstochastic control theory takes the best of both worlds, simultaneously granting robustness to adversarial noise and allowing the use of optimistic behavior when appropriate, while providing computationally efficient methods.
@@ 1.6 The Need for a New Theory
Consider the problem of flying a drone from source to destination subject to unknown weather conditions. The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight. However, the wind conditions, rain and other uncertainties are a different story. Certainly the wind is not an i.i.d. Gaussian random variable! Optimal control theory, which assumes this zero-mean noise, is therefore overly optimistic and inaccurate for our problem.
@@ 1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, i.e. arbitrary or even adversarial perturbations, the optimal policy, i.e. a decision making rule mapping observations to control inputs, is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can completely be predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games – regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$. Then the controller observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$. Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
@@
At this point the reader may wonder why we should compare to a reference class, as opposed to all possible policies. There are two main answers to this question:
@@
    First, the best policy in hindsight out of all possible policies may be very complicated to describe and reason about. This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
@@
    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs. the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed, as we discuss later in the book.
@@ 1.6.2 A New Family of Algorithms
As a representative example, consider the Gradient Perturbation Controller described in Algorithm 1. The algorithm maintains matrices that are used to create the control as a linear function of the past perturbations, as per line 3. These parameters change over time according to a gradient-based update rule, where the gradient is calculated as the derivative of a counterfactual cost function with respect to the parametrization, as per line 5.
@@
The version given here assumes that the time-invariant dynamics $f$ is known. The GPC computes a control which is a linear function of past perturbations. It modifies the linear parameters according to the loss function via an iterative gradient method. In Chapter 7, we prove that this algorithm attains sublinear regret bounds for a large class of time-varying and time-invariant dynamical systems. In later chapters we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
@@ 1.7 Bibliographic Remarks
Control theory is a mature and active discipline with both deep mathematical foundations and a rich history of applications. In terms of application, self-regulating feedback mechanisms that limit the flow of water, much like the modern float valve, were reportedly used in ancient Greece over two thousand years ago. The first mathematical analysis of such systems appeared in a paper [Max68] by James Clerk Maxwell on the topic of governors, shortly after Maxwell had published his treatise on the unification of electricity and magnetism. For an enlightening historical account, see [FCZI03].
For a historical account on the development of the PID controller see [Ben93]. The theoretical and practical properties of this controller are spelled out in detail in [ÅH95].
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers an extensive coverage of topics from the viewpoint of dynamic programming. The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory. Rockafellar [Roc87] proposed the use of convex cost functions for modeling the cost in a linear dynamical system to incorporate constraints on the state and control, and suggested computational difficulties with this approach. The text [SL91] gives a thorough introduction to adaptive control theory.
Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19]. A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21]. Notable amongst them are logarithmic regret algorithms [AHS19, FS20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CHK+18], and nonstochastic control with partial observability [SSH20].
The theory builds heavily on recent developments in decision making and online learning, notably the framework of online convex optimization [Haz16], and especially its extensions to learning with memory [AHM13].
@@
In this chapter we return to the general problem of control from definition 1.2, i.e. choosing a sequence of controls to minimize a sequence of cost functions. Before moving to algorithms for control, it is important to reason about what kind of solutions are even possible.
For this reason we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior. This will help us understand the kinds of observations and control objectives that are feasible and natural in real applications.

### v5

# version 0.9

@@
[Uncaptioned image] Introduction to Online Control
version 0.9
@@
We are in a golden age for control and decision making. A proliferation of new applications including self-driving vehicles, humanoid robots, and artificially intelligent drones open a new set of challenges for control theory to address. Hazan and Singh have written the definitive book on the New Control Theory– non-stochastic control. The phrase “a paradigm shift” has become cliche from overuse, but here it is truly well deserved; the authors have revisited the foundations by focusing on building controllers that perform nearly as well as if they knew future disturbances in advance, rather than relying on probabilistic or worst-case models. The non-stochastic control approach has extended one of the most profound ideas in mathematics of the 20th century, online (no-regret) learning, to master sequential decision making with continuous actions. This leads to high performance in benign environments and resilience in adversarial ones. The book, authored by pioneers in the field, presents both foundational concepts and the latest research, making it an invaluable resource.
Drew Bagnell
Yann LeCun
To my family: Dana, Hadar, Yoav, Oded and Deluca,
—EH
To my parents– A & V.
—KS
© [2025] Elad Hazan and Karan Singh
@@ Acknowledgments
This book was a journey that started in Princeton at the Google DeepMind Lab in the summer of 2018. At that time Sham Kakade visited the group, which consisted of the authors together with Naman Agarwal, Brian Bullins, Xinyi Chen, Cyril Zhang, and Yi Zhang, henceforth the gang. The gang worked on various aspects of optimization, reinforcement learning, and dynamical systems. The first paper, in which online nonstochastic control was basically invented, started in room 401 in the computer science building at Princeton University. We remember and cherish the exact moment along with the mathematical writings in white chalk on that blackboard.
@@
The following colleagues and students were particularly helpful, making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna, and Jonathan Pillow (and many more we wish to announce very soon).
@@
Linear Algebra, Geometry and Calculus
@@
$A\succeq 0$ $A$ is positive semi-definite, $\forall v\ ,\ v^{\top}Av\geq 0$
$A\succeq B$ $A-B$ is positive semi-definite
@@ 1.1 What is This Book About?
Control theory is the engineering science of manipulating physical systems to achieve the desired functionality. Its history is centuries old and spans rich areas from the mathematical sciences. A brief list includes differential equations, operator theory and linear algebra, functional analysis, harmonic analysis, and more. However, it mostly does not concern modern algorithmic theory or computational complexity from computer science.
The advance of the deep learning revolution has brought to light the power and robustness of algorithms based on gradient methods. The following quote is attributed to deep learning pioneer Yann Lecun during a Barbados workshop on deep learning theory circa February 2020.
@@
This is a deep and difficult question that we fall short of answering. However, we do propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
@@ 1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century. One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism. A major source of efficiency losses of the earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily cause engine failure. The centrifugal governor was created to solve this problem.
Figure 1.1: A centrifugal governor.
The centrifugal governor consists of a shaft connected to two metal balls that are allowed to rotate. The increased shaft speed causes the balls to spin faster and the resulting centrifugal force causes the balls to lift. The balls are in turn attached to a pin that controls the amount of steam entering the engine, so as the balls spin faster, the pin lifts and limits the amount of fuel that can enter the system. However, if the balls spin slower, the pin is lowered, allowing more steam to enter the system. This provides a natural, direct, and proportional control to the system.
The impact of the centrifugal governor on the steam engine, and in turn on the industrial revolution, illustrates the importance of the field of control in engineering and the sciences. Since the 18th century, our world has changed quite a bit, and the importance of control has only grown.
The challenges of controlling systems have moved from the steam engine to the design of aircraft, robotics, and autonomous vehicles. The importance of designing efficient and robust methods remains as crucial as in the industrial revolution.
@@ 1.3 Formalization and Examples of a Control Problem
With the physical example of the centrifugal governor in mind, we give a mathematical definition of a control problem. The first step is to formally define a dynamical system in the following way.
@@
A dynamical system evolving from an initial state $\mathbf{x}_{1}$ is given by the following equations:
@@
The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given the current state $\mathbf{x}_{t}$, the input control $\mathbf{u}_{t}$, and the perturbation $\mathbf{w}_{t}$. The subscript indicates the relevant quantities at time step $t$.
Given a dynamical system, there are several possible objectives that we can formulate. We give a generic formulation below that captures many common control formulations, notably the optimal control problem and the robust control problem. These will be addressed in much more detail later in this book. In this problem formulation, the goal is to minimize the long-term horizon cost given by a sequence of cost functions.
@@
At this point, many aspects of the generic problem have not been specified. These include:
@@
Different theories of control differ by the answers they give to the above questions, and these in turn give rise to different methods.
Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem. We can then reason about the different questions and motivate the nonstochastc control setting.
@@ 1.3.1 Example: Control of a Medical Ventilator
The ventilator must take into account the structure of the lung to determine the optimal pressure to induce. Such structural factors include compliance, or the change in lung volume per unit pressure, and resistance, or the change in pressure per unit flow.
@@
where $p_{t}$ is the observed pressure, $v_{t}$ is the volume of the lung, $u_{t}$ is the input flow (control) and $\Delta_{t}$ is the unit of time resolution of the device. $C_{0}$, $C_{1}$, and $C_{2}$ are constants that depend on the particular system/lung-ventilator pair.
The goal of ventilator control is to regulate the pressure according to some predetermined pattern. As a dynamical system, we can write $v_{t+1}=f(v_{t},u_{t},w_{t})$, $p_{t}=g(v_{t})$, and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$. The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
@@ 1.4 Simple Control Algorithms
The previous section detailed the importance of control problems, specifically through the example of a critical care medical device. In this section, we detail methods for controlling a medical ventilator, or indeed any physical device that can be modeled as a dynamical system.
The following methods are simplistic, but still widely used. For example, the PID algorithm is still used in most medical ventilators today. In later sections, we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
@@ 1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a prespecified dynamic range $[x_{\min},x_{\max}]$ (which can vary over time) and where the scalar control input $u_{t}$ must at all times be within the interval $[u_{\min},u_{\max}]$, which is assumed to contain zero for simplicity. A bang-bang controller chooses the control input at time $t$ as
@@
A bang-bang controller saturates its inputs to extreme values in an attempt to confine the state of the system to a desirable range. For example, in the context of mechanical ventilation, one could choose $[p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$. In such a case, a bang-bang controller would allow oxygen to enter at the highest possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
@@ 1.4.2 The PID Controller
A special subfamily of linear controllers is that of linear controllers that optimize on a predefined basis of historical observations. More precisely,
@@
The PID controller, named after the nature of the three components that make up it, specifies three coefficients $\alpha,\beta,\gamma$ to generate a control signal. In that sense, it is extremely sparse, which from a learning-theoretic perspective is a good indication of generalization.
In fact, the PID controller is the method of choice for numerous engineering applications, including control of medical ventilators.
@@ 1.5 Classical Theory: Optimal and Robust Control
Classical theories of control differ within the generic control problem formulation 1.2 primarily in one significant aspect: the perturbation model. A rough divide between the two main theories – optimal control and robust control – is as follows.
Optimal control postulates a probabilistic model for the perturbations, which are typically i.i.d. draws from a Gaussian distribution. However, other types of distribution can also be considered. As such, the control problem can be re-phrased in a more precise stochastic mathematical program,
@@
In contrast, robust control theory plans for the worst-case noise from a given set of constraints, denoted $\mathcal{K}$. Thus, the robust control problem can be rephrased as the following mathematical program,
@@
The two formulations have been debated for decades, with pros and cons for each side. Optimal control focuses on the average case and often yields simpler formulations for optimization, resulting in more efficient methods. Robust control, on the other hand, allows for more difficult noise models and some amount of model misspecification. However, it is too pessimistic for many natural problems.
Nonstochastic control theory takes the best of both worlds, simultaneously granting robustness to adversarial noise and allowing the use of optimistic behavior when appropriate while providing computationally efficient methods.
@@ 1.6 The Need for a New Theory
Consider the problem of flying a drone from the source to the destination subject to unknown weather conditions. The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight. However, the wind conditions, rain, and other uncertainties are a different story. Certainly, the wind is not an i.i.d. Gaussian random variable! Optimal control theory, which assumes this zero-mean noise, is therefore overly optimistic and inaccurate for our problem.
@@ 1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, i.e. arbitrary or even adversarial perturbations, the optimal policy, i.e. a decision making rule mapping observations to control inputs, is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can be completely predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games, namely regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$. The controller then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$. Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
@@
At this point, the reader may wonder why we should compare to a reference class, as opposed to all possible policies. There are two main answers to this question:
@@
    First, the best policy in hindsight out of all possible policies may be very complicated to describe and to reason about. This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
@@
    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret versus the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed, as we discuss later in the book.
@@ 1.6.2 A New Family of Algorithms
As a representative example, consider the Gradient Perturbation Controller described in Algorithm 1. The algorithm maintains matrices that are used to create the control as a linear function of past perturbations, according to line 3. These parameters change over time according to a gradient-based update rule, where the gradient is calculated as the derivative of a counterfactual cost function with respect to the parametrization, according to line 5.
@@
The version given here assumes that the time-invariant dynamics $f$ is known. The GPC computes a control that is a linear function of past perturbations. It modifies the linear parameters according to the loss function via an iterative gradient method. In Chapter 7, we prove that this algorithm attains sublinear regret bounds for a large class of time-varying and time-invariant dynamical systems. In later chapters, we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
@@ 1.7 Bibliographic Remarks
Control theory is a mature and active discipline with both deep mathematical foundations and a rich history of applications. In terms of application, self-regulating feedback mechanisms that limit the flow of water, much like the modern float valve, were reportedly used in ancient Greece over two thousand years ago. The first mathematical analysis of such systems appeared in a paper by James Clerk Maxwell on the topic of governors [Max68], shortly after Maxwell had published his treatise on the unification of electricity and magnetism. For an enlightening historical account, see [FCZI03].
For a historical account on the development of the PID controller, see [Ben93]. The theoretical and practical properties of this controller are described in detail in [ÅH95].
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers extensive coverage of topics from the point of view of dynamic programming. The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory. Rockafellar [Roc87] proposed the use of convex cost functions to model the cost in a linear dynamical system to incorporate constraints on state and control and suggested computational difficulties with this approach. The text [SL91] gives a thorough introduction to adaptive control theory.
Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19]. A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21]. Notable among them are logarithmic regret algorithms [AHS19, FS20, Sim20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CH21], and nonstochastic control with partial observability [SSH20].
The theory is heavily based on recent developments in decision making and online learning, notably the framework of online convex optimization [Haz16], and especially its extensions to learning with memory [AHM13].
@@
In this chapter, we return to the general problem of control from Definition 1.2, that is, choosing a sequence of controls to minimize a sequence of cost functions. Before moving on to algorithms for control, it is important to reason about what kind of solution is even possible.
For this reason, we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior. This will help us understand the kinds of observation models and control objectives that are feasible and natural in real applications.